In a survey, 3/16 of people said that they preferred to use self-service gas stations. 5/8 said that they preferred not to pump their own gas. The remaining 75 respondents said that they had no clear preference. How many people preferred self service? Can you please explain to me how to work this problem? Thank you
Answer by sudhanshu_kmr(1152)


Let total number of person = x
Number of person that preferred to use self-service = 3x/16
number of persons that not preferred to use self-service = 5x/8
and number of remaining persons that had no clear preference = 75
these three groups form total number of persons...
3x/16 + 5x/8 + 75 = x
=> 13x/16 + 75 = x
=> 75 = x - 13x/16
=> 75 = 3x/16
so, x = 400
Number of person that preferred to use self-service = 3/16 * 400 = 75